OpenStudy (anonymous):

For which \(a\) does the following series converge? \[\sum_{n=2}^{\infty}\frac{1}{n(\ln n)^a}\]

OpenStudy (anonymous):

For which \(a\) does the following series converge? \[\sum_{n=2}^{\infty}\frac{1}{n(\ln n)^a}\]

OpenStudy (anonymous):

i think for "a" from (0,+inf)

OpenStudy (anonymous):

does not converge for a=1

OpenStudy (anonymous):

I can tell you that it converges for a>= 2 but I don't know the reason yet

OpenStudy (anonymous):

We might have to use comparison test

OpenStudy (zarkon):

integral test

OpenStudy (zarkon):

have you figured it out yet :)

OpenStudy (zarkon):

foods ready..be back

OpenStudy (zarkon):

a>1

OpenStudy (anonymous):

\[a_n=\left(\frac{1}{n*\text{Log}[n]^2}\right)\] \[b_n=\frac{1}{n^2}\] \[\frac{a_n}{b_n}\] as n goes to infinity \[\frac{\left(\frac{1}{n*\text{Log}[n]^2}\right)}{\frac{1}{n^2}}\] \[n\rightarrow \infty \]\[\frac{n}{\text{Log}[n]^2}\] =0 According to rules two of limit comparison test , it converge http://archives.math.utk.edu/visual.calculus/6/series.9/index.html

OpenStudy (zarkon):

is should converge for a=2 since it converges for all a>1

OpenStudy (anonymous):

a>1, bingo. Thanks!

OpenStudy (zarkon):

@imranmeah91: that limit above is not zero

OpenStudy (anonymous):

what do you mean?

OpenStudy (zarkon):

\[\lim_{n\to\infty}\frac{n}{\text{Log}[n]^2}\neq0\]

OpenStudy (anonymous):

nvm, got it